# Enhanced sensitivity to trace 238U impurity of sapphire via coincidence neutron activation analysis

### v1

@@ I Introduction
The development of the new analysis technique discussed in this paper was motivated by the low-radioactivity material requirements of large time projection chambers (TPCs) being used in ongoing [7, 8, 9, 10, 11, 12] and planned rare event searches [13, 14, 15], with specific application to the development of the next-generation double beta decay search nEXO [16]. Filled with different noble elements, their ability to reconstruct event energies, tracks, and locations gives them a powerful background recognition capability, and the use of a fluid target offers scalability to ton-size devices. Achieving a uniform electric drift field across the TPC requires the installation of internal electrode structures, which need to be mechanically supported and electrically isolated from other components. For low-rate, low-energy experiments, the support material, as all other components, needs to be as radiopure as possible.
@@ II Measurement setup
A $2.5482(3)\,\text{g}$ sample of sapphire, received in 2015 from Saint-Gobain Crystals, New Hampshire (now Luxium Solutions), was obtained from a larger crystal that was itself saw cut, ground, and annealed at Saint-Gobain. The sample was subsequently cut to the largest size that would fit inside of an activation vial by the University of Massachusetts Amherst glass shop. It was then washed in an ultrasonic bath of acetone and then ethanol, for ten minutes each, before shipping to Alabama, were it was again cleaned and packaged for activation at MIT.
@@
The sapphire sample was activated at the MIT research reactor on May 22, 2024 for a duration of $8.28000(28)\,\text{h}$. This irradiation was immediately followed by the activation of two standards, consisting of NIST reference material SRM 1633b (fly ash) with masses of $0.4059(3)\,\text{g}$ and $0.5083(3)\,\text{g}$. The standards were activated for only 187(1) s to avoid creating dangerously high activities. The 2PH1 sample insertion facility was used for all samples as it offers the highest thermal neutron flux. The sapphire and one fly ash sample were sent to TUNL for $\upgamma\upgamma\text{ coincidence counting}$. To determine the neutron flux, and to ensure analysis consistency between the two facilities, the other fly ash sample was sent to the University of Alabama (UA) for single $\upgamma\text{ counting}$. In all cases, the end of activation established time zero for the following analyses. The samples were delivered to both analysis labs on May 24, 2024.
@@
The TUNL $\upgamma\upgamma\text{ coincidence-counting}$ setup consist of two ORTEC coaxial high-purity germanium (HPGe) detectors, nominally 88 mm in diameter and 49 mm in length. The detectors are surrounded by passive shielding consisting of 2 cm of Cu and 15 cm of Pb. A 3D-printed sample holder is utilized to keep the two HPGe detectors collinear and to center the sample between the two detectors. This setup was previously used for studies of $\upbeta\upbeta$ decay to excited final states [22, 23] at the Kimballton Underground Research Facility. It was relocated to ground level at Duke University prior to the present work.
@@
The HPGe detector efficiencies were determined using commercial $\gamma$-ray calibration sources from Eckert & Ziegler ($^{60}\text{Co}$, and a scatterless source containing $^{54}\text{Mn}$, $^{57}\text{Co}$, $^{65}\text{Zn}$, $^{88}\text{Y}$, $^{109}\text{Cd}$, $^{113}\text{Sn}$, $^{134}\text{Cs}$ and $^{137}\text{Cs}$, $^{139}\text{Ce}$, and $^{241}\text{Am}$). These sources were chosen as they span a wide range of energies, from 60–1836 keV, and emit coincident $\gamma$ rays ($^{60}\text{Co}$, $^{88}\text{Y}$, and $^{134}\text{Cs}$). Source calibration was performed before and after sample counting, in both the 1 cm and the 5 cm counting geometry. Data taken with these activity-calibrated sources served to tune and verify the validity of the detector response model, described in Sec. III.
@@ III Detector response modeling
Figure 2: Simulated detector geometry. Detectors separated by 2 cm. The sapphire is shown inside a sample counting bottle; the two calibration source geometries (gray) are shown inside a source holder with different thickness spacers. The sample holder, that sets the separation between the detectors, is not modeled.
@@
The physical characteristics of the HPGe detectors—in particular, their dead layer, the geometry of their germanium crystals, and the separation between the detectors—were known with limited precision. The scatterless source with mixed isotopes, with a broad spectrum of activity and coincident $\gamma$ rays of $^{88}\text{Y}$, was used in the simulation to adjust the parameters related to these unknowns. A dead layer for both detectors of about 1.6 mm, a true separation of 2.6 cm, and a source offset by 1 mm, were found to make the source derived detector efficiency more compatible with that derived from the simulation.
The residual discrepancy between the efficiency derived from calibration source data and that derived from the simulation is shown in Fig. 3.
@@
Figure 3: Ratio of calibration source derived detection efficiency to simulation derived detection efficiency for prominent $\gamma$ rays from the mixed source (see text for a list of nuclides) and from the $^{60}\text{Co}$ source (open squares). Ge-A in orange; Ge-B in blue. Data points are averages over several calibration runs with RMS error. Top are for single $\gamma$ rays. Bottom includes coincident pairs of the most prominent $\gamma$ rays from $^{60}\text{Co}$, $^{88}\text{Y}$, and $^{134}\text{Cs}$. For coincidences, the points in orange are representative of analysis of the same data as the blue.
@@
Only scatterless source data that was not used to tune the simulation (independent detector calibration data) is shown. Several $^{60}\text{Co}$ source runs, with more complex source geometry when compared to the scatterless source, are also shown. The agreement between the source data and simulation validates a 10% systematic uncertainty we apply to single $\upgamma\text{-ray}$ efficiencies derived from this simulation.
@@
The coincident detection efficiency derived from the source data is compared to that derived from the simulation in Fig. 3, bottom. Pairs of the two most prominent $\upgamma\text{-ray}$ transitions—of $^{60}\text{Co}$, $^{88}\text{Y}$, and $^{134}\text{Cs}$, seperately—were analyzed requiring coincidence with each other (blue and orange in the figure). These analysis differ slightly, depending on which detector is designated as the tagging detector; the same signal events are analyzed, only the background events may differ. The lack of variation between the permutations in Fig. 3, bottom, indicates that this analysis is insensitive to the choice of which detector serves as the tagging detector. In general, agreement is observed that validates a 10% systematic uncertainty on coincident $\upgamma\text{-ray}$ efficiency derived from the simulation.
@@
In addition to general errors derived from evaluating sample efficiency with this simulation, there are several associated specifically with simulating the sapphire sample. The dimensions and density of the sapphire sample, and its position and orientation within the counting vial, were all varied in the simulation. The position of the sample ($\pm 4.5\,\text{mm}$) within the counting vial (14.5 mm diameter) altered the simulated sapphire single $\upgamma\text{-ray}$ detection efficiency within a $\pm 7\%$ (relative) range. The orientation of the sample, tilting towards one detector or the other, increased the simulated sapphire coincident $\upgamma\text{-ray}$ detection efficiency by up to 3% (relative). This uncertainty is subdominant to, and to some extent correlated with, the detector response model uncertainty as measured with the calibration sources.
@@ IV Data analysis and results
NAA determines the summed contributions of surface and bulk impurities. The measurement reported here states contaminant concentrations normalized by mass. To suppress contributions from surface impurities, the sample was carefully cleaned by employing a series of sequential ultrasonic baths in solutions of 2% micro-90 detergent, 50% $\rm HNO_{3}$, 50% HCl, and water using a procedure developed at Pacific Northwest National Laboratory (PNNL). The need for careful and thorough chemical cleaning of the surface was evident from early sapphire measurements. Earlier assays of sapphire, that were not effectively cleaned, measured $^{238}\text{U}$ concentrations in the parts-per-billion range. After more robust cleaning, employing the PNNL method, replicate samples were at levels determined in this work. PNNL measured the leachates from the cleanings using inductively coupled plasma mass spectrometry to ensure iterative sonication leachate levels reached a baseline before being submitted for NAA.
After cleaning, samples were packaged in similarly cleaned polyethylene (PE) bottles, which are required to separate them from the environment inside the research reactor. Sample handling was performed in a class 500 clean room at UA. Only ultrapure chemicals were used. After their return to the respective counting labs, the samples were removed from their co-activated PE bottle and transferred into a clean counting bottle. Post activation, sample cleanliness is no longer an issue, as activities due to short-lived nuclides are not naturally occurring. The post activation handling, instead, focuses on safety and containment of the activity.
@@ IV.1 Neutron flux
[…]
@@
[…]
A point-wise systematic error of 2.8% has been applied to the extrapolated fly ash activity to achieve a $\chi^{2}=0.99$ for three degrees of freedom between the observed and modeled activity when extracting these fluxes. Average neutron capture cross sections were taken from the JENDL-4.0 library [27]. A detailed description of an equivalent neutron flux determination can be found in Ref. [26].
@@ IV.2 Data analysis
Single $\upgamma\text{-ray}$ spectra, not requiring a time coincidence, were analyzed separately for detectors Ge-A and Ge-B. For the $\upgamma\upgamma$ coincidence analysis, events in one detector were selected with a coincidence condition applied to the other; requiring the presence of an event within a predefined energy region and coincidence time window in the second detector. The energy regions were determined from the tabulated $\gamma$ cascade energies of the activation products. They were defined as the total absorption peak $\pm 3\sigma$ (energy resolution).
@@
The time dependence of the counting rates was modeled with an exponential function, fixing the half-life to that of the nuclide identified through the choice of full absorption peak. The analysis has high selectivity because it is double differential in energy and time. Extracted counting rates, extrapolated to the end of activation, along with the detector response model derived efficiency and measured neutron flux, were then used to quantify elemental concentrations in the fly ash and sapphire samples.
Because this study uses a novel approach and a detection system differing from previous studies, we decided to also use an alternative approach to the conversion of counting rates into element concentrations. This approach is less reliant on the detector response modeling. Both the standard fly ash material and the sapphire were exposed to the same neutron environment and counted with the same detector setup, with similar geometry. The ratio of counting rates, eliminates the dependence on measured neutron flux and capture cross sections. The elemental concentration in the sample, $c_{\text{s}}$, was calculated from the ratio of sample to fly ash count rates, extrapolated to the end-of-activation $r_{\text{s}}(0)\big{/}r_{\text{a}}(0)$, as
@@ IV.3 Quality checks
Figure 4: Time dependence of the observed count rates for four coincident photopeaks associated with $^{24}\text{Na}$, $^{46}\text{Sc}$, $^{99}\text{Mo}$, and $^{181}\text{Hf}$ present in the sapphire sample. An exponential decay model, both with fixed and free half-life parameter, are shown after a least squares fit to the normalization of the data points.
Because this was the first use of the described method, care was taken to validate results. The following tests were performed to validate the consistency of the method with previous measurements:
@@
    In 2017, elemental concentrations of Sc, Cr, Co, Mo and Hf were determined at UA by activating and single $\upgamma\text{ counting}$ the same sapphire sample. The 1PH1 sample insertion facility was used. It has a different neutron energy spectrum compared to 2PH1. TUNL single-$\upgamma$ and $\upgamma\upgamma\text{ coincidence counting}$ results were compared with the 2017 UA data.
@@
    For Cr, the elemental concentration evaluated via the measured neutron flux, capture cross sections and detector response model-derived efficiency was compared with that resulting from the fly ash method and single $\upgamma\text{ counting}$.
@@
    The data is modeled with an exponential decay; both with the half-life fixed to that of the expected isotope (taken from [29, 30, 31, 32]), and with the half-life as a free parameter. The average ratio of the expected-to-free half-lives is $0.9774\pm 0.0015$. The rate uncertainty due to sample positioning is not included in these least square fits of this model to the data and, therefore, not reflected in the uncertainty on the extracted half-lives. Allowing the normalization of the various distinct groups of data points, each group representing data taken with the sample in precisely the same position, reproduces the 7% sapphire sample position uncertainty introduced in Sec. III. Given the small statistics, in the number of distinct sample positions, the affect of this uncertainty cannot be well established with this data; whereas the statistical uncertainty on the individual extracted rates are well established.

### v3

@@ I Introduction
The development of the new analysis technique, discussed in this paper, was motivated by the low-radioactivity material requirements of large time projection chambers (TPCs) being used in ongoing [8, 7, 2, 30, 27, 12] and planned rare event searches [3, 1, 5], with specific application to the development of the next-generation double beta decay search nEXO [6]. Filled with different noble elements, their ability to reconstruct event energies, tracks, and locations gives them a powerful background recognition capability, and the use of a fluid target offers scalability to ton-size devices. Achieving a uniform electric drift field across the TPC requires the installation of internal electrode structures, which need to be mechanically supported and electrically isolated from other components. For low-rate, low-energy experiments, the support material, as all other components, needs to be as radiopure as possible.
@@ II Measurement setup
Figure 2: The sapphire sample is shown (middle) together with its PE counting vial (left) and snap lid (right).
@@
A $2.5482(3)\,\text{g}$ sample of sapphire, received in 2015 from Saint-Gobain Crystals, New Hampshire (now Luxium Solutions), was obtained from a larger crystal that was itself saw cut, ground, and annealed at Saint-Gobain. The sample designation is G3. The sample was subsequently cut to the largest size that would fit inside of an activation vial by the University of Massachusetts Amherst glass shop. The sapphire sample has approximately cuboid shape with side dimensions of $\rm(3.8-4.0)\;mm\times(3.6-4.1)\;mm\times 34.5\;mm$, as shown in Fig. 2. Prior to shipping to the University of Alabama (UA), it was washed in an ultrasonic bath of acetone and then ethanol, for ten minutes each.
NAA determines the summed contributions of surface and bulk impurities. The measurement reported here states contaminant concentrations normalized by mass. To suppress contributions from surface impurities, the sample was carefully cleaned at Pacific Northwest National Laboratory (PNNL) before activation at MIT. After that the sample was only handled in clean room environments or a fume hood.
The precision cleaning procedure developed at PNNL employed a series of sequential ultrasonic baths in ultrapure solutions of 2% micro-90 detergent, 50% $\rm HNO_{3}$, 50% HCl, and water. The need for careful and thorough chemical cleaning of the surface was evident from a 2016 measurement of Saint-Gobain sapphire sample A1 (same material as G3 but different surface treatment). This assay, also by NAA, did not use an effective cleaning protocol and, as a result, measured $^{238}\text{U}$ concentrations in the parts-per-billion range. PNNL measured the leachates from the cleanings using inductively coupled plasma mass spectrometry. This ensured that iterative sonication leachate levels had reached a baseline before submitting the cleaned sample for NAA. A 2017 NAA re-assay of the PNNL-cleaned sapphire sample G3 yielded only upper limits for $\rm{}^{232}Th$ and $\rm{}^{238}U$ in the ppt-range.
After cleaning, samples were packaged in similarly cleaned polyethylene (PE) bottles, which are required to separate the sapphire from the environment inside the research reactor. Pre-activation sample handling was performed in a class 500 clean room at UA.
After sample return to the respective counting labs, the sample was removed from its co-activated PE bottle and transferred into a clean cylindrical PE counting bottle with $\rm 11.7\;mm$ diameter and $\rm 54.5\;mm$ height. Post activation, sample cleanliness is no longer an issue, as activities due to short-lived nuclides are not naturally occurring. Typical sample-induced counting rates are of order kHz, much higher than detector background. The post activation handling, instead, focuses on safety and containment of the activity.
@@
The sapphire sample was activated at the MIT research reactor on May 22, 2024 for a duration of $8.28000(28)\,\text{h}$. This irradiation was immediately followed by the activation of two standards, consisting of NIST reference material SRM 1633b (fly ash) with masses of $0.4059(3)\,\text{g}$ and $0.5083(3)\,\text{g}$. The standards were activated for only 187(1) s to avoid creating dangerously high activities. The 2PH1 sample insertion facility was used for all samples as it offers the highest thermal neutron flux. The sapphire and one fly ash sample were sent to TUNL for $\upgamma\upgamma\text{ coincidence counting}$. To determine the neutron flux, and to ensure analysis consistency between the two facilities, the other fly ash sample was sent to UA for single $\upgamma\text{ counting}$. In all cases, the end of activation established time zero for the following analyses. The fly ash samples were delivered to both analysis labs on May 24, 2024. After concluding counting of the sapphire sample at TUNL, it was sent to UA to assay long-lived activities and establish analysis consistency. The sample was received at UA on August 20, 2025, 90 days after the end of activation.
@@
The TUNL $\upgamma\upgamma\text{ coincidence-counting}$ setup consists of two ORTEC $p$-type coaxial high-purity germanium (HPGe) detectors. The detectors have nominally identical geometries: 88 mm crystal diameter, 49 mm crystal length, and a 2.54 mm thick nickel-plated magnesium window. The two detectors are herein referred to as Ge-A and Ge-B and have a resolution at 122 keV of 1.8 and 1.1 keV full width at half maximum (FWHM), respectively. The Ge-A and Ge-B resolution at 1332 keV is 2.7 and 2.0 keV FWHM, respectively. The detectors are surrounded by passive shielding consisting of 2 cm of Cu and 15 cm of Pb. A 3D-printed sample holder is utilized to keep the two HPGe detectors collinear and to center the sample between the two detectors. This setup was previously used for studies of $\upbeta\upbeta$ decay to excited final states [24, 22] at the Kimballton Underground Research Facility. It was relocated to ground level at Duke University prior to the present work.
@@
The HPGe detector efficiencies were determined using commercial $\gamma$-ray calibration sources from Eckert & Ziegler. $^{60}\text{Co}$ radioactivity was used in form of a button source. A so-called “scatterless” source contains $^{54}\text{Mn}$, $^{57}\text{Co}$, $^{65}\text{Zn}$, $^{88}\text{Y}$, $^{109}\text{Cd}$, $^{113}\text{Sn}$, $^{134}\text{Cs}$, $^{137}\text{Cs}$, $^{139}\text{Ce}$, and $^{241}\text{Am}$, sandwiched between two thin aluminized Mylar sheets. These sources were chosen as they span a wide range of energies, from 60–1836 keV, and emit coincident $\gamma$ rays ($^{60}\text{Co}$, $^{88}\text{Y}$, and $^{134}\text{Cs}$). Source calibrations were performed before and after sample counting, in both the 1 cm and the 5 cm counting geometry. Data taken with these activity-calibrated sources served to tune and verify the validity of the detector response model, described in Sec. III.
@@ III Detector response modeling
Figure 3: Simulated detector geometry. Detectors separated by 2 cm. The sapphire is shown inside a sample counting bottle. Because of the size mismatch between sapphire sample and counting bottle, the location and orientation of the sample were not well defined. Simulations were performed for various orientations. The top panel shows the sample leaning diagonally inside the bottle; its most likely orientation. The two calibration source geometries (gray) are shown inside a source holder with different thickness spacers. The sample holder, which sets the separation between the detectors, is not modeled.
@@
The physical characteristics of the HPGe detectors—in particular, their dead layer, the geometry of their germanium crystals, and the separation between the detectors—were known with limited precision. The scatterless source with mixed isotopes, with a broad spectrum of activity, was used in the simulation to adjust the parameters related to these unknowns. A dead layer for both detectors of about 1.6 mm, a true separation of 2.6 cm, and a source offset by 1 mm, were found to make the source-derived detector efficiency more compatible with that derived from the simulation.
@@
Figure 4: Ratio of calibration source-derived detection efficiency to simulation-derived detection efficiency for prominent $\gamma$ rays from the scatterless source (see text for a list of nuclides) and from the $^{60}\text{Co}$ source (open squares). Ge-A in orange; Ge-B in blue. Data points are averages over several calibration runs with their RMS errors. Top are for single $\gamma$ rays. Bottom includes coincident pairs of the most prominent $\gamma$ rays from $^{60}\text{Co}$, $^{88}\text{Y}$, and $^{134}\text{Cs}$. Both permutations of the coincidence condition are shown in the lower figure: like pairs of orange points and like pairs of blue points are not independent measurements; the analysis does not depend on the choice of tagging detector.
@@
The residual discrepancy between the efficiency derived from calibration source data and that derived from the simulation is shown in Fig. 4. Only scatterless source data that was not used to tune the simulation (independent detector calibration data) is shown. Several $^{60}\text{Co}$ source runs, with more complex source geometry when compared to the scatterless source, are also shown. The agreement between the source data and simulation validates a 10% systematic uncertainty we apply to single $\upgamma\text{-ray}$ efficiencies derived from this simulation.
@@
The coincident detection efficiency derived from the source data is compared to that derived from the simulation in Fig. 4, bottom. Pairs of the two most prominent $\upgamma\text{-ray}$ transitions—of $^{60}\text{Co}$, $^{88}\text{Y}$, and $^{134}\text{Cs}$, separately —were analyzed, requiring coincidence with each other (blue and orange in the figure). These analysis differ slightly, depending on which detector is designated as the tagging detector; the same signal events are analyzed, only the background events may differ. The lack of variation between the permutations in Fig. 4, bottom, indicates that this analysis is insensitive to the choice of which detector serves as the tagging detector. In general, agreement is observed that validates a 10% systematic uncertainty on coincident $\upgamma\text{-ray}$ efficiency derived from the simulation.
@@
In addition to general errors arising from evaluating sample efficiency with this simulation, there are several specifically associated with simulating the sapphire sample. The dimensions and density of the sapphire sample, and its position and orientation within the counting vial, were all varied in the simulation. The dimensions and density of the sample are well known, resulting in small contributions to the overall uncertainty. The uncertainty in the position of the sample dominated the total error on the single $\upgamma\text{-ray}$ counting efficiency. In contrast, the orientation of the sample dominated the error on the coincident $\upgamma\text{-ray}$ efficiency. The position of the sample within the counting vial altered the simulated sapphire single $\upgamma\text{-ray}$ detection efficiency within a $\pm 7\%$ (relative) range for either detector. The orientation of the sample, tilting towards one detector or the other, increased the simulated sapphire coincident $\upgamma\text{-ray}$ detection efficiency by up to 3% (relative). This uncertainty is subdominant to, and to some extent correlated with, the detector response model uncertainty as measured with the calibration sources.
@@ IV.1 Neutron flux
The UA fly ash counting data was used to determine that the thermal $\Phi_{\text{th}}$, epithermal $\Phi_{\text{et}}$, and fast neutron flux $\Phi_{\text{f}}$, for this MIT reactor activation, were
@@
$\Phi_{\text{et}}$ $=$ $\mathrm{(1.50\pm 0.12)\cdot 10^{13}\,\frac{neutrons}{cm^{2}\cdot s}}$
$\Phi_{\text{f}}$ $=$ $\mathrm{(7.39\pm 0.88)\cdot 10^{12}\;\frac{neutrons}{cm^{2}\cdot s}}$
A point-wise systematic error of 2.8% has been added in quadrature to the extrapolated fly ash activities to achieve $\chi^{2}=0.99$ for three degrees of freedom between the observed and modeled activities. Average neutron capture cross sections were taken from the JENDL-4.0 library [32]. A detailed description of an equivalent neutron flux determination can be found in Ref. [25].
@@ IV.2 Data analysis
Single $\upgamma\text{-ray}$ spectra, not requiring a time coincidence, were analyzed separately for detectors Ge-A and Ge-B. Both detectors provide independent event triggers. The single-event distributions contain no linkage between the two detectors and represent different photon populations. They are, therefore, independent of each other. These independent data sets are combined by means of a weighted average, to reduce uncertainty. The simulations of the two detectors were tuned separately using radioactive source data. We, therefore, assume that their simulation uncertainties are independent of each other. The systematic uncertainty was included in the weights during averaging.
For the $\upgamma\upgamma$ coincidence analysis, events in one detector were selected with a coincidence condition applied to the other; requiring the presence of an event within a predefined energy region and coincidence time window in the second, the tagging detector. The energy regions were determined from the tabulated $\gamma$ cascade energies of the activation products. They were defined as the total absorption peak $\pm 3\sigma$ (energy resolution). The full absorption peak of photon 1 was fit in data collected with Ge-A, with Ge-B acting as the tagger, required to fully absorb coincident photon 2. Requiring photon 2 to be absorbed by Ge-A and tagging photon 1 in Ge-B selects different photons. These two data sets are statistically independent and were combined by averaging. However, the same photons are selected when merely exchanging the roles of the fit with that of the tagging detector. Rates resulting from this permutation of roles were not included in the average.
@@
The time dependence of the counting rates was modeled with an exponential function, fixing the half-life to that of the nuclide identified through the choice of full absorption peak. The analysis has high selectivity because it is double differential in energy and time. Extracted counting rates, extrapolated to the end of activation, along with the detector response model-derived efficiency and measured neutron flux, were then used to quantify elemental concentrations in the fly ash and sapphire samples.
Because this study uses a novel approach and a detection system differing from previous studies, we decided to also use an alternative approach to the conversion of counting rates into element concentrations. This approach is less reliant on the detector response modeling. Both the standard fly ash material and the sapphire were exposed to the same neutron environment and counted with the same detector setup, with similar geometry. The ratio of counting rates eliminates the dependence on measured neutron flux and capture cross sections. The elemental concentration in the sample, $c_{\text{s}}$, was calculated from the ratio of sample to fly ash count rates, extrapolated to the end-of-activation $r_{\text{s}}(0)\big/r_{\text{a}}(0)$, as
@@ IV.3 Quality checks
Figure 5: Time dependence of the observed count rates for four coincident photopeaks associated with $^{24}\text{Na}$, $^{46}\text{Sc}$, $^{99}\text{Mo}$, and $^{181}\text{Hf}$ present in the sapphire sample. An exponential decay model, both with fixed (indicated by the isotope) and free half-life parameter (shown with uncertainty), are shown after a least squares fit to the normalization of the data points. The lines conform to the color coding of the text.
Because this was the first use of the described method, care was taken to validate results. The following tests were performed to ensure the consistency of the method with previous measurements:
@@
    In 2017, elemental concentrations of Sc, Cr, Co, Mo and Hf were determined at UA by activating and single $\upgamma\text{ counting}$ the same Saint-Gobain G3 sapphire sample. In 2017 the 1PH1 sample insertion facility was used. It has a different neutron energy spectrum compared to 2PH1  used for this analysis. TUNL single-$\upgamma$ and $\upgamma\upgamma\text{ coincidence counting}$ results were compared with the 2017 UA data.
@@
    For Cr in Saint-Gobain G3 sapphire, the elemental concentration evaluated via the measured neutron flux, capture cross sections and detector response model-derived efficiency was compared with that resulting from the fly ash method and single $\upgamma\text{ counting}$ at TUNL.
@@
    The data is modeled with an exponential decay; both with the half-life fixed to that of the expected isotope (taken from [31, 17, 38, 37]), and with the half-life as a free parameter. The average ratio of the expected-to-free half-lives is $0.9774\pm 0.0015$. The rate uncertainty due to sample positioning is not included in the least square fits of this model to the data and, therefore, not reflected in the uncertainty of the extracted half-lives. Allowing the normalization of the various distinct groups of data points to float, each group representing data taken with the sample in precisely the same position, improves the agreement with the data, and the agreement with the expected half-life. However, this model was not generally applied to all the data since the individual normalization of the separate positions must be averaged to obtain the initial activity of the sample; the variation is accounted for by the application of the more robust position derived uncertainty from section III. We consider the observation that measured and tabulated half-lives agreed, on average, within 2.3% as evidence that the elemental species can be properly identified by this type of decay time analysis.